OpenStudy (anonymous):

WILL MEDAL.... Solve each equation. Check your answer: 1.6m=1.28 Thanks

OpenStudy (anonymous):

1.6m = 1.28 1.6m/1.6 = 1.28/1.6 m = 0.8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob